OpenStudy (anonymous):

int x^2 sin(2x) dx can someone please show me steps

OpenStudy (anonymous):

got it myself thanks everyone

OpenStudy (anonymous):

u= x^2 , dv= sin2x => du = 2x , v = -1/2* cos2x and

OpenStudy (experimentx):

use integration by parts ... use x^2 as u,

OpenStudy (experimentx):

i guess you have to use integration by parts twice in a row ..

OpenStudy (anonymous):